OpenStudy (anonymous):

Solve the compound inequality. 6>-2x+5 or 10< or equal to -3x+5

OpenStudy (anonymous):

\[[1> - 2x] or [5 \le -3x\]] \[[-1/2 < x]or [-5/3 \ge x]\] \[x < -1/2\] OR \[x \ge -5/3\]

OpenStudy (anonymous):

x>-1/2 the first one

OpenStudy (anonymous):

x<=-5/3

OpenStudy (anonymous):

2nd one